OpenStudy (anonymous):

2/3(6c-18)+3/4(8c+32)=-18

OpenStudy (anonymous):

\[\frac{2}{3} (-18+6 c)+\frac{3}{4} (32+8 c)=-18 \]Multiply each side by 12 and simplify.\[12\left(\frac{2}{3} (6 c-18)+\frac{3}{4} (8 c+32)\right)=12(-18) \]\[144+120 c=-216 \]\[c=-3 \]

OpenStudy (anonymous):

appreciate it man

OpenStudy (anonymous):

You're welcome.